OpenStudy (anonymous):

There are 2 cell phone plans. Plan 1 gives a free phone, $1 per day used fee and then 8 cents per minute. Plan 2 gives a free phone as well, has a monthly fee of $10.00 and 18 cents per minute. If someone decides to use their phone for 15 days a month only, at what # of minutes do the plans cost the same?

OpenStudy (anonymous):

15+8x=10+18x 5=10x 1/2=x half minute huh?

OpenStudy (anonymous):

that doesnt check though. i got 11.5 = 14

OpenStudy (anonymous):

hmm

OpenStudy (mathmate):

15+0.08x=10+0.18x (0.18-0.08)x = 15-10 x=(15-10)/(0.18-0.08)=50 So at 50 minutes (over 15 days), the two plans cost the same.